### 10 Mar 2026 v2

# On the generalized circular projected Cauchy distribution

Abstract
Tsagris and Alzeley, (2025) proposed the generalized circular projected Cauchy distribution, whose special case is the wrapped Cauchy distribution. In this paper we first derive the relationship with the wrapped Cauchy distribution, and then we attempt to characerize the distribution. We provide a non-analytical formula for the mean resultant length, and analytical form for the cumulative probability function and the the entropy. We propose a log-likelihood ratio test for the equality of two angular means, without assuming equality of thew concentration parameters, and briefly mention the addition of circular and simplicial predictors. Simulation studies illustrate a) the performance of the log-likelihood ratio test when one falsely assumes that the true distribution is the wrapped Cauchy distribution, and b) the empirical rate of covergence of the regression coefficients. Using a real data analysis example we correct a mistake in the regression setting.
Keywords: Circular data; mean resultant length, hypothesis test
MSC: 62H11, 62H10
1 Introduction
Directional data refers to multivariate data with a unit norm, whose sample space can be expressed as:
[…]
where $\left|\left|.\right|\right|$ denotes the Euclidean norm. Circular data, when $d=2$, lie on a circle. Circular data are encountered in various disciplines, such as political sciences (Gill and Hangartner,, 2010), criminology (Shirota and Gelfand,, 2017), biology (Landler et al.,, 2018), ecology (Horne et al.,, 2007), and astronomy (Soler et al.,, 2019), to name but a few.
A large class of distributions has been proposed, see Tsagris and Alzeley, (2025) for a short list. In this paper we focus on the generalized circular projected Cauchy distribution (GCPC) proposed by Tsagris and Alzeley, (2025). Specifically we show the relationship with the wrapped Cauchy (WC) distribution and further propose a log-likelihood ratio test for the equality of two independent angular means.
The next section briefly presents the GCPC distribution, derives the relationship with the WC distribution and proposes the log-likelihood ratio test. Then, simulation studies follow and the conclusions that close the paper.
2 The GCPC distribution
[…]
$f({\bf y})=\int_{0}^{\infty}r^{d-1}f(r{\bf y})dr.$ (1)
The probability density function of the bivariate Cauchy distribution, with some location vector $\boldsymbol{\mu}$ and scatter matrix $\boldsymbol{\Sigma}$, is given by
$f({\bf x})=\frac{1}{2\pi|\boldsymbol{\Sigma}|^{1/2}}\left[1+\left({\bf x}-\boldsymbol{\mu}\right)^{\top}\boldsymbol{\Sigma}^{-1}\left({\bf x}-\boldsymbol{\mu}\right)\right]^{-3/2}.$ (2)
By substituting (2) into (1) and evaluating the integral, Tsagris and Alzeley, (2025) derived the termed the Circular Projected Cauchy (CPC) distribution.
$f({\bf y})$ $=$ $\int_{0}^{\infty}\frac{r}{2\pi|\boldsymbol{\Sigma}|^{1/2}}\left(1+r^{2}{\bf y}^{\top}\boldsymbol{\Sigma}^{-1}{\bf y}-2r{\bf y}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\mu}+\boldsymbol{\mu}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\mu}\right)^{-3/2}dr$ (3)
$=$ $\frac{1}{2\pi|\boldsymbol{\Sigma}|^{1/2}\left(B\sqrt{\Gamma^{2}+1}-A\sqrt{B}\right)},$
where
$A={\bf y}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\mu},\ B={\bf y}^{\top}\boldsymbol{\Sigma}^{-1}{\bf y},\ \text{and}\ \Gamma^{2}=\boldsymbol{\mu}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\mu}.$
It is important to note that ${\bf y}\in\mathbb{S}^{1}$, while $\bm{\mu}\in\mathbb{R}^{2}$.
(Tsagris and Alzeley,, 2025) relaxed this strict assumption by employing one of the conditions imposed in Paine et al., (2018), that is, $\boldsymbol{\Sigma\mu}=\boldsymbol{\mu}$, but not $|\boldsymbol{\Sigma}|=1$. This condition implies that the one eigenvector $\boldsymbol{\xi}_{2}$ of $\boldsymbol{\Sigma}$ is the normalised location vector $\boldsymbol{\xi}_{2}=\left(\mu_{1},\mu_{2}\right)^{\top}/\gamma$, while the other eigenvector can be defined up to the sign as $\boldsymbol{\xi}_{1}=\left(-\mu_{2},\mu_{1}\right)^{\top}/\gamma$ or $\boldsymbol{\xi}_{1}=\left(\mu_{2},-\mu_{1}\right)^{\top}/\gamma$. The eigenvalue corresponding to the location vector is equal to 1, while the other eigenvalue is equal to $\lambda$, hence $|\boldsymbol{\Sigma}|=\lambda>0$ and the inverse of the scatter matrix is given by
$\boldsymbol{\Sigma}^{-1}=\frac{1}{\gamma^{2}}\left(\begin{array}[]{cc}\mu_{1}^{2}+\mu_{2}^{2}/\lambda&\mu_{1}\mu_{2}\left(1-1/\lambda\right)\\ \mu_{1}\mu_{2}\left(1-1/\lambda\right)&\mu_{2}^{2}+\mu_{1}^{2}/\lambda\end{array}\right)=\boldsymbol{\xi}_{1}\boldsymbol{\xi}_{1}^{\top}/\lambda+\boldsymbol{\xi}_{2}\boldsymbol{\xi}_{2}^{\top}.$ (6)
Thus, (3) becomes
$f({\bf y})=\frac{1}{2\pi\lambda^{1/2}\left(B\sqrt{\gamma^{2}+1}-a\sqrt{B}\right)}.$ (7)
Utilising (6) and after some calculations, the density in (7) may also be expressed in polar coordinates by
$f(\theta)$ $=$ $\frac{\left(2\pi\lambda^{1/2}\right)^{-1}}{\left[\left(\cos^{2}(\theta-\omega)+\frac{\sin^{2}(\theta-\omega)}{\lambda}\right)\sqrt{\gamma^{2}+1}-\gamma\cos(\theta-\omega)\sqrt{\cos^{2}(\theta-\omega)+\frac{\sin^{2}(\theta-\omega)}{\lambda}}\right]}$ (8)
$=$ $\frac{1}{2\pi\lambda^{1/2}\left(b\sqrt{\gamma^{2}+1}-a\sqrt{b}\right)},$
where $a=\gamma\cos\phi,\ b=\cos^{2}\phi+\frac{\sin^{2}\phi}{\lambda},\ \phi=\theta-\omega$, with $\lambda>0$, $\gamma\geq 0$ and $\omega\in[-\pi,\pi]$.
The GCPC distribution exhibits symmetry with respect to $\omega$ such that $f(\theta-\omega)=f(\omega-\theta)$, and is thus symmetrical about $\omega+\pi$. The density of the GCPC may be re-written as
$f(\theta)=\frac{1}{2\pi\sqrt{\lambda}\sqrt{b}D}=\frac{1}{2\pi\sqrt{1+\left(\lambda-1\right)\cos^{2}{\phi}}\cdot D},$ (9)
where
$D=c\sqrt{b}-\gamma\cos\phi\ \text{and}\ c=\sqrt{\gamma^{2}+1}.$
It is important to note that if $\lambda=1$, the GCPC distribution reduces to the circular independent projected Cauchy (CIPC) distribution (Tsagris et al.,, 2025)
$f_{CIPC}({\bf\theta})=\dfrac{1}{2{\pi}\left(\sqrt{\gamma^{2}+1}-\gamma\cos{\left(\theta-\omega\right)}\right)},$ (10)
which is the WC distribution (Mardia and Jupp,, 2000, pg. 51) with a different parameterization
$f_{WC}(\theta)=\frac{1-\delta^{2}}{2\pi\left[1+\delta^{2}-2\delta\cos\left(\theta-\omega\right)\right]},$ (11)
where $\gamma=\frac{2\delta}{1-\delta^{2}}$ or, conversely, $\delta=(\sqrt{\gamma^{2}+1}-1)/\gamma$ (Tsagris and Alzeley,, 2025).
2.1 Relationship with the CIPC distribution
Theorem 2.1.
[…]
$\arctan 2(y,x)=\left\{\begin{array}[]{ll}\arctan(y/x)&\text{if}\ x>0,\\ \arctan(y/x)+\pi&\text{if}\ x<0\ \text{and}\ y\geq 0,\\ \arctan(y/x)-\pi&\text{if}\ x<0\ \text{and}\ y<0,\\ \pi/2&\text{if}\ x=0\ \text{and}\ y>0,\\ -\pi/2&\text{if}\ x=0\ \text{and}\ y<0,\\ \text{undefined}&\text{if}\ x=y=0.\end{array}\right$. (18)
Proof.
The proof is straightfoward by application of the change-of-variables formula to obtain
$f_{\Psi}(\psi)=f_{\Theta}(\theta)\cdot\left|\frac{d\theta}{d\psi}\right|=\frac{1}{2\pi\sqrt{\lambda}\,\sqrt{b}\,D}\cdot\sqrt{\lambda}\cdot b=\frac{1}{2\pi\!\left(\sqrt{\gamma^{2}+1}-\gamma\cos\psi\right)}.$
2.2 Probabilities
Following Tsagris and Alzeley, (2025), instead of the cumulative probability function we derive the probability included within a given interval $(a,b)$, and provide an alternative formula.
$P(a\leq\theta\leq b)=\int_{a}^{b}f_{GCPC}(t)dt=\int_{-\pi}^{b}f_{GCPC}(t)dt-\int_{-\pi}^{a}f_{GCPC}(t)dt=F_{GCPC}(b)-F_{GCPC}(a).$
Applying Theorem 2.1 the formula above becomes:
$P(a\leq\theta\leq b)$ $=$ $\int_{-\pi}^{\psi(b)}f_{WC}(\psi)\,d\psi-\int_{-\pi}^{\psi(a)}f_{WC}(\psi)\,d\psi=F_{WC}(\psi(b))-F_{WC}(\psi(a))$
$=$ $\left[\frac{1}{2}+\frac{1}{\pi}\arctan\left(\frac{1+\delta}{1-\delta}\tan\frac{\psi(b)}{2}\right)\right]-\left[\frac{1}{2}+\frac{1}{\pi}\arctan\!\left(\frac{1+\delta}{1-\delta}\tan\frac{\psi(a)}{2}\right)\right]$
$=$ $\frac{1}{\pi}\left[\arctan\!\left(\frac{1+\delta}{1-\delta}\tan\frac{\psi(b)}{2}\right)-\arctan\left(\frac{1+\delta}{1-\delta}\tan\frac{\psi(a)}{2}\right)\right]$
where $\psi(t)=\arctan 2\left(\sin(t-\omega),\,\sqrt{\lambda}\cos(t-\omega)\right)$.
2.3 Mean resultant length
The mean resultant length is defined as $\rho=E\left[\cos{\left(\theta-\omega\right)}\right]$. Based on this, one may compute the circular variance as $1-\rho$, and the circular standard deviation as $\left(-2\log{\rho}\right)^{1/2}$. Using Theorem 2.1 we may express this expectation as
$\rho=\frac{1-\delta^{2}}{\pi}\int_{0}^{\pi}\frac{\cos\psi}{\sqrt{\cos^{2}\psi+\lambda\sin^{2}\psi}}\cdot\frac{1}{1+\delta^{2}-2\delta\cos\psi}\,d\psi.$
After some algebra we end up with the following representation
$\rho=\frac{2(1+\delta)}{\pi(1-\delta)\sqrt{\lambda}}\;\Pi\!\left(\left(\frac{1+\delta}{1-\delta}\right)^{2}\frac{1}{\lambda}-1\;\Bigg|\;1-\frac{1}{\lambda^{2}}\right),$
where $\Pi(a|b)$ is the complete elliptic integral of the third kind (Ward,, 1960). We still see that there is no closed form solution for $\rho$, unless $\lambda=1$, and this applies to higher trigonometric moments. Figure 1(a) visualizes the values of $\rho$ for a grid of values of $\gamma$ and $\lambda$. We observe that $\rho$ increases with increasing $\gamma$ and decreases with increasing $\lambda$ values.
2.4 Entropy
Theorem 2.2.
The entropy of the GCPC distribution is given by
$H_{GCPC}(\omega,\gamma,\lambda)=\log{\left\{\frac{8\pi\sqrt{\lambda}\,(1-\delta^{2})}{\bigl[\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}\bigr]^{2}}\right\}},$
where $\delta$ was defined earlier.
Proof.
The entropy is defined as $H=-\int_{-\pi}^{\pi}f(\theta)\log f(\theta)\,d\theta$. Taking the logarithm of (9), we have: $\log f_{GCPC}=-\log(2\pi)-\frac{1}{2}\log\lambda-\frac{1}{2}\log b-\log D$. Using the change of variables from Theorem 2.1, we have $f_{GCPC}(\theta)\,d\theta=f_{WC}(\psi)\,d\psi$ and
$\log D=\frac{1}{2}\log b+\log(c-\gamma\cos\psi)\ \ \text{and}\ \ b\left[\phi(\psi)\right]=\frac{1}{1+(\lambda-1)\sin^{2}\psi},$
where $c-\gamma\cos\psi=(1+\delta^{2}-2\delta\cos\psi)/(1-\delta^{2})$.
By changing variables and substituting the result into the entropy we obtain:
$H_{GCPC}=\log(2\pi)+\frac{1}{2}\log\lambda+E_{WC}\left(\log b\right)+\int_{-\pi}^{\pi}f_{WC}(\psi)\log(c-\gamma\cos\psi)\,d\psi.$
The last integral equals $H_{WC}-\log{\left(2\pi\right)}$, thus
$H_{GCPC}=H_{WC}+\frac{1}{2}\log\lambda-E_{WC}\left[\log(1+(\lambda-1)\sin^{2}\psi)\right].$ (19)
We know that the entropy of the WC distribution is $H_{WC}=\log{\left[2\pi\left(1-\delta^{2}\right)\right]}$. Let us now write $1+(\lambda-1)\sin^{2}\psi=\frac{1+\lambda}{2}(1-\alpha\cos 2\psi)$ with $\alpha=(\lambda-1)/(\lambda+1)$. Then, $1-\alpha\cos 2\psi=(1+r_{1}^{2}-2r_{1}\cos 2\psi)/(1+r_{1}^{2})$ where
$r_{1}=\frac{1-\sqrt{1-\alpha^{2}}}{\alpha}=\frac{\sqrt{\lambda}-1}{\sqrt{\lambda}+1}.$
Using the Fourier identity $\log(1+\delta^{2}-2\delta\cos\psi)=-2\sum_{n=1}^{\infty}\frac{\delta^{n}}{n}\cos(n\psi)$ we get $E_{\mathrm{WC}}[\cos(n\psi)]=\delta^{n}$. Applying the same Fourier identity at frequency $2n$, with $E_{WC}\left[\cos(2n\psi)\right]=\delta^{2n}$:
$E_{WC}\left[\log(1-\alpha\cos 2\psi)\right]=-\log(1+r_{1}^{2})+2\log(1-r_{1}\delta^{2}).$
Using $\log\frac{1+\lambda}{2}-\log(1+r_{1}^{2})=2\log\frac{\sqrt{\lambda}+1}{2}$ and $1-r_{1}\delta^{2}=\frac{\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}}{\sqrt{\lambda}+1}$ we obtain
$E_{WC}\left[\log(1+(\lambda-1)\sin^{2}\psi)\right]=2\log\frac{\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}}{2}.$
Substituting the above result into (19) we obtain
$H_{GCPC}(\omega,\gamma,\lambda)=\log{\left\{\frac{8\pi\sqrt{\lambda}\,(1-\delta^{2})}{\bigl[\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}\bigr]^{2}}\right\}}.$
Note that if $\lambda=1$: $H_{GCPC}=\log(2\pi(1-\delta^{2}))=H_{WC}$, and if $\delta=0$: $H_{GCPC}=\log(8\pi\sqrt{\lambda}/(\sqrt{\lambda}+1)^{2})\leq\log(2\pi)$. Figure 1(b) visualizes the values of $H$ for a grid of values of $\gamma$ and $\lambda$. We observe that, in cotrast to $\rho$, $H$ increases with increasing $\lambda$ and decreases with increasing $\gamma$ values.
Refer to caption Refer to caption
(a) $\rho$ (b) $H$
Figure 1: Values of the mean resultant length, $\rho$ and entropy, $H$, as a function of $\gamma$ and $\lambda$.
2.5 Kullback-Leibler divergence between GCPC and CIPC
[…]
$KL(GCPC\|CIPC)$ $=$ $-\frac{1}{2}\log\lambda+2\log\!\left(\frac{\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}}{2}\right)-\log(1-\delta^{2})$
$+E_{WC}\!\left[\log\!\left(\sqrt{\gamma^{2}+1}-\frac{\gamma\cos\psi}{\sqrt{1+(\lambda-1)\sin^{2}\psi}}\right)\right],$
where the expectation is taken with respect to $\psi\sim WC(0,\delta)$.
2.6 Hypothesis testing and confidence intervals
2.6.1 Hypothesis test for the location parameter of one sample
In order to test whether the location parameter equals some pre-specified value we will employ the log-likelihood ratio test. Under the $H_{0}$, $\omega=\omega_{0}$, and so we need to maximize the restricted log-likelihood, $\ell_{0}$, with respect to the other two parameters, $\gamma$ and $\lambda$, and estimate their values, $\tilde{gamma}$ and $\tilde{\lambda}$. Under the $H_{1}$, $\omega\neq\omega_{0}$, and we maximize the unrestricted log-likelihood, $\ell_{1}$, to obtain $\hat{\omega}$, $\hat{\gamma}$ and $\hat{\lambda}$. Under regularity conditions, $\Lambda=2\left[\ell_{1}\left(\hat{\omega},\hat{\omega},\hat{\lambda}\right)-\ell_{0}\left(\omega_{0},\tilde{\gamma},\tilde{\lambda}\right)\right]\mathrel{\dot{\sim}}\chi^{2}_{1}$.
2.6.2 Confidence interval for a location parameter
Using the log-likelihood ratio test, we can construct asymptotic confidence intervals for the true location parameter $\omega$ of a sample by searching for the pair of values for which it holds that (Cox and Hinkley,, 1979): $\ell\left(\omega,\hat{\gamma},\hat{\lambda}\right)>\ell\left(\hat{\omega},\hat{\gamma},\hat{\lambda}\right)-\frac{\chi^{2}_{1}}{2}$.
2.6.3 Equality of two location parameters
To test the equality of two location parameters, without assuming equality of the parameters $\gamma$ and $\lambda$, we will perform a log-likelihood ratio test [^1] The function to perform the log-likelihood ratio test is available in the R package Directional (Tsagris et al.,, 2026). following (Tsagris et al.,, 2025). Assume we have circular observations from two independent samples, $(\theta_{11},\ldots,\theta_{1{n_{1}}})$ and $(\theta_{11},\ldots,\theta_{1{n_{2}}})$, where $n_{1}$ and $n_{2}$ denote the sample sizes of the two samples.
Under $H_{0}$, $\omega_{1}=\omega_{2}=\tilde{\omega}$, and by using Eq. (8), the log-likelihood is written as
$\ell_{0}\left(\tilde{\omega},\tilde{\lambda}_{1},\tilde{\lambda}_{2},\tilde{\gamma}_{1},\tilde{\gamma}_{2}\right)$ $=$ $-(n_{1}+n_{2})\log(2\pi)-\frac{n_{1}}{2}\log(\tilde{\lambda}_{1})-\frac{n_{2}}{2}\log(\tilde{\lambda}_{2})$
$-\sum_{i=1}^{n_{1}}\log{\left(\tilde{b}_{1i}\sqrt{\tilde{\gamma}_{1}^{2}+1}-\tilde{a}_{1i}\sqrt{\tilde{b}_{1i}}\right)}-\sum_{i=1}^{n_{2}}\log{\left(\tilde{b}_{2i}\sqrt{\tilde{\gamma}_{2}^{2}+1}-\tilde{a}_{2i}\sqrt{\tilde{b}_{2i}}\right)},$
where $\tilde{a}_{ji}=e^{\tilde{\gamma}_{j}}\cos{\left(\theta_{ji}-\tilde{\omega}\right)}$ and $\tilde{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\tilde{\omega}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\tilde{\omega}\right)}{\tilde{\lambda}_{j}}$, for $j=1,2$.
Under $H_{1}$, $\omega_{1}\neq\omega_{2}$, and the log-likelihood is written as
$\ell_{1}\left(\hat{\omega},\hat{\lambda}_{1},\hat{\lambda}_{2},\hat{\gamma}_{1},\hat{\gamma}_{2}\right)$ $=$ $-(n_{1}+n_{2})\log(2\pi)-\frac{n_{1}}{2}\log(\hat{\lambda}_{1})-\frac{n_{2}}{2}\log(\hat{\lambda}_{2})$
$-\sum_{i=1}^{n_{1}}\log{\left(\hat{b}_{1i}\sqrt{\gamma_{1}^{2}+1}-\hat{a}_{1i}\sqrt{\hat{b}_{1i}}\right)}-\sum_{i=1}^{n_{2}}\log{\left(\hat{b}_{2i}\sqrt{\hat{\gamma}_{2}^{2}+1}-\hat{a}_{2i}\sqrt{\hat{b}_{2i}}\right)},$
where $\hat{a}_{ji}=e^{\hat{\gamma}_{j}}\cos{\left(\theta_{ji}-\hat{\omega}_{j}\right)}$ and $\hat{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\hat{\omega}_{j}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\hat{\omega}_{j}\right)}{\hat{\lambda}_{j}}$, for $j=1,2$.
Under regularity conditions, $\Lambda=2\left[\ell_{1}\left(\hat{\omega},\hat{\lambda}_{1},\hat{\lambda}_{2},\hat{\gamma}_{1},\hat{\gamma}_{2}\right)-\ell_{0}\left(\tilde{\omega},\tilde{\lambda}_{1},\tilde{\lambda}_{2},\tilde{\gamma}_{1},\tilde{\gamma}_{2}\right)\right]\mathrel{\dot{\sim}}\chi^{2}_{1}$.
2.7 Regression modelling revisited
Tsagris and Alzeley, (2025) defined the GCPC regression by considering the bivariate representation (Euclidean coordinates) of the circular data as opposed to their univariate nature (Kato et al.,, 2008). This is akin to the approach employed in the SPML model, as detailed by Presnell et al., (1998) and allows for varying concentration parameter ($\lambda$). The log-likelihood of the GCPC regression model is written as
$\ell_{GCPC}$ $=$ $-\sum_{i=1}^{n}\log{\left(\bm{y}_{i}^{\top}\bm{\Sigma}^{-1}_{i}\left(\bm{B}\right)\bm{y}_{i}\sqrt{\bm{\mu}_{i}^{\top}\bm{\mu}_{i}+1}-\bm{y}_{i}^{\top}\bm{\mu}_{i}\sqrt{{\bf y}_{i}^{\top}\bm{\Sigma}^{-1}_{i}\left(\bm{B}\right)\bm{y}_{i}}\right)}$ (20)
$-\frac{n}{2}\log(\lambda)-n\log{(2\pi)},$
where
$\bm{y}_{i}^{\top}\bm{\Sigma}^{-1}_{i}\left(\bm{B}\right)\bm{y}_{i}$ $=$ $\xi_{1i}\left(\bm{B}\right)\xi_{1i}\left(\bm{B}\right)^{\top}/\lambda+\xi_{2i}\left(\bm{B}\right)\xi_{2i}\left(\bm{B}\right)^{\top}$
$=$ $y_{1i}^{2}\left(\xi_{1i}^{2}/\lambda+\xi_{2i}^{2}\right)+y_{2i}^{2}\left(\xi_{2i}^{2}/\lambda+\xi_{1i}^{2}\right)+2y_{1i}y_{2i}\xi_{1i}\xi_{2i}\left(1-1/\lambda\right)$
and
$\xi_{ji}=\frac{\bm{\beta}_{j}^{\top}x_{i}}{\|\bm{\beta}_{j}^{\top}x_{i}\|}=\frac{\mu_{ji}}{\|\bm{\mu}_{i}\|}=\frac{\mu_{ji}}{\gamma_{i}}\ \ \text{for}\ \ j=1,2,$ (21)
where $\bm{x}_{i}$ denoting the $i$-th observation of the covariate vector.
To maximize the log-likelihood of the GCPC regression model (20) suggested to use multiple starting points, but we propose a faster option. We begin with initial regression coefficients produced by the SPML regression model of Presnell et al., (1998) and then estimate the $\lambda$ parameter using Brent’s algorithm [^2] This is available in R via the buit-in optimize() function. (Brent,, 2013). Then use these as starting points in R’s built-in optim() function.
Tsagris and Alzeley, (2025) did not consider the case of circular-circular regreression, where a covariate $X$ is circular as well. But, this case is easily adopted in the above case scenario by transforming the circular covariate to Euclidean by using $\bm{z}_{i}=\left(\cos{x_{i}},\sin{x_{i}}\right)^{\top}$. In case of multiple circular covariates, all of them are transformed into their Euclidean coordinates and the same link function (21) is used. Simplicial predictors (compositional data) are straighforward to add by using two options. The first is to apply the additive log-ratio transformation (Aitchison,, 2003), so that the regression coefficients sum to zero and have a nice (derivative based) interpretation. In case of zero values present the logarithmic transformation breaks down, so the alternative approach is to transform them via the $\alpha$–transformation (Tsagris et al.,, 2011). The second approach bypasses the problem of zero values but lacks interpretation of the estimated regression coefficients and requires to choose the value of the $\alpha$ parameter. The choice of $\alpha$ may be overcome by choosing a small value of $\alpha$ (say 0.01) to resemble the isometric log-ratio transformation (Egozcue et al.,, 2003).
An extra extension would be to link the $\lambda$ parameter to the covariates as well, but that would increase the computational cost. Another evident approach to make the model non-linear would be to use splines in the covariates as well.
3 Simulation studies
We performed simulation studies to examine the performance of the two sample means test and also the effect of the $\lambda$ parameter on the performance of the log-likelihood test.
3.1 One location parameter
We selected six sample sizes, and generated values from the $GCPC\left(2,3,2\right)$. For each sample size we computed the type I error (assuming $\alpha=0.05$) based on 1,000 replicates. Table 1 presents the results, where we can see that the test slightly overestimates the nominal type I error, but not significantly.
Table 1: Estimated type I error of the log-likelihood ratio test for one location parameter assuming the GCPC distribution.
Sample size 30 50 70 100 150 200
Type I error 0.065 0.059 0.066 0.062 0.067 0.063
3.2 Two location parameters
We compared the test when both samples come from the GCPC distribution and when assuming that they come from the CIPC distribution ($\lambda=1$). Six pairs of unequal sample sizes were used, where the true angular mean was $\omega=2$, but the $\gamma$ and $\lambda$ parameters differed. The parameters for the smaller sample were equal to $\gamma=4$ and $\lambda=1$, while for the larger sample they were equal to $\gamma=2$ and $\lambda=3$. Circular data were generated from two GCPC distributions with the specified parameters and the two log-likelihood ratio tests were performed. This process was repeated 1,000 times and the proportion of rejection of the $H_{0}$ (at the $\alpha=0.05$ level) served as an estimate of the type I error. Table 2 presents the results. The GCPC based log-likelihood ratio always attained the correct size, whereas the CIPC based log-likelihood ratio test overestimated the size of the test.
Table 2: Estimated type I error of the log-likelihood ratio test for the equality of two location parameters, assuming the GCPC distribution and assuming the CIPC distribution.
Sample size GCPC CIPC
(30, 50) 0.054 0.098
(30, 70) 0.056 0.094
(30, 100) 0.060 0.102
(50, 70) 0.048 0.100
(50, 100) 0.059 0.101
(70, 100) 0.061 0.092
3.3 Empirical asymptotics for the regression model
We estimated empirically the order of convergence of the regression coefficients, with a) one circular predictor, b) one continuous predictor and c) the combination of both. The grid for the sample sizes considered ranged from 100 up to 5,000 at an increasing step of 100. For every sample size $n$, data were generated with some pre-specified coefficients and the regression model was fitted.
For the circular predictor, the model was
$\left(y_{1},\ y_{2}\right)=\left(1,\ \cos{u},\ \sin{u}\right)\left(\begin{array}[]{cc}\beta_{01}&\beta_{02}\\ \beta_{11}&\beta_{12}\\ \beta_{21}&\beta_{22}\end{array}\right),$
where $\bm{u}$ was generated from a von Mises distribution with mean equal to 2 and concentration parameter equal to 5. For the continuous predictor, the model was
$\left(y_{1},\ y_{2}\right)=\left(1,\ x\right)\left(\begin{array}[]{cc}\beta_{01}&\beta_{02}\\ \beta_{11}&\beta_{12}\\ \end{array}\right),$
where $\bm{x}$ was generated from an exponential distribution with mean equal to 0.5. Table 3 shows the matrix with the ground truth regression coefficients $\bm{B}$ used. For all cases, the $\lambda$ values of the GCPC distribution used to generate the circular responses were equal to 0.5 and 5.
The discrepancy between the observed and the estimated regression coefficients was measured using the squared Frobenius norm. This process was repeated 50 times and the average squared Frobenius norm ($\bar{F}^{2}$) was stored. Then, we estimated the coefficient of the following regression model:
$\log{\bar{F}^{2}}\sim a+b\log{n}.$
The estimated $b$ served as the empirical rate of convergence of the regression coefficients.
Table 3 contains the matrix $\bm{B}$ of the regression parameters used in the simulation study. Figure 2 visualizes the rates when $\lambda=0.5$ and $\lambda=5$. For all cases the estimated slope is very close to 1, indicating that the convergence rate of the estimated regression coefficients is $n^{-1}$.
Table 3: Matrix $\bm{B}$ of the regression coefficients used in the simulations. The first two columns correspond to the coefficients of the model with the circular predictor, the second two to the model with the continuous predictor and the last two to the model with both of them. In the last case $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$ refer to the circular predictor and $\bm{\beta}_{3}$ to the continous predictor.
Predictor Circular Continuous Circular and continuous
$\cos(y)$ $\sin(y)$ $\cos(y)$ $\sin(y)$ $\cos(y)$ $\sin(y)$
$\bm{\beta}_{0}$ 1.874 2.550 1.641 1.949 1.859 3.508
$\bm{\beta}_{1}$ -1.473 -2.023 -0.101 -0.119 -1.500 -2.914
$\bm{\beta}_{2}$ -1.157 -1.554 -0.960 -1.855
$\bm{\beta}_{3}$ -0.246 -0.181
Circular covariate
Refer to caption Refer to caption
(a) $\lambda=0.5$ (b) $\lambda=5$
Continuous covariate
Refer to caption Refer to caption
(c) $\lambda=0.5$ (d) $\lambda=5$
Circular and continuous covariate
Refer to caption Refer to caption
(e) $\lambda=0.5$ (f) $\lambda=5$
Figure 2: Average squared Frobenius norm $\bar{F}^{2}$ versus sample size, both in log-scale for the circular, the continuous and the combination of both.
4 Real data analysis
We used the same dataset used in Tsagris and Alzeley, (2025), the speed.wind2 data that is available in the R package NPCirc (Oliveira et al.,, 2014). This dataset consists of 199 hourly observations of wind direction and wind speed in winter season (from November to February) from 2003 until 2012 in the Atlantic coast of Galicia (Spain).
4.1 Maximum likelihood estimation
Table 4 presents the estimated parameters of the CIPC and GCPC models. Application of the log-likelihood ratio test to discriminate between the two models (Tsagris and Alzeley,, 2025) clearly favors the GCPC model. Figure 3(a) visualizes the data, it is the circular plot with the mean directions showed with arrows. Evidently the data look bimodal and the CIPC has chosen one of the two modes, whereas the GCPC has chosen a direction that ”cuts” the data into two. The density plot on Figure 3(b) shows that the GCPC density has a similar pattern to the one produced by the kernel density estimate, in contrast to the CIPC distribution that has failed to fit the data. This is more evident in the circular density plot (Figure 3(c)). The CIPC distribution has a circular shape, whereas the GCPC distribution has an elliptical shape, capturing the shape of the data more accurately. Finally, Figure 3(d) visualizes the 95% confidence interval for the mean direction.
Table 4: Estimated parameters of the CIPC and GCPC models applied to the wind direction data.
Model $\hat{\omega}$ $\hat{\gamma}$ $\hat{\lambda}$ Log-likelihood
CIPC 0.603 0.203 -363.930
GCPC 5.587 0.050 4.21 -337.739
Refer to caption Refer to caption
(a) Circular plot (b) Density plot
Refer to caption Refer to caption
(c) Circular density plot (d) 95% confidence interval for the mean direction
Figure 3: (a) Circular plot of the speed direction data. (b) Density plot of the kernel density estimate, and the fitted CIPC and GCPC distributions. (c) Circular density plot of the two fitted distributions. The green line is the estimated mean direction based on the CIPC, whereas the blue lines is the estimated mean direction based on the GCPC model.
4.2 Regression
We spotted a mistake in the regression modelling of Tsagris and Alzeley, (2025), where the continuous predictor is the speed of the wind. Table 5 shows the regression coefficients computed in Tsagris and Alzeley, (2025) and the correct estimates. Note that this time the sign of the slope coefficients matches that of the CIPC and of the spherically projected multivariate linear regression models models. The $\lambda$ coefficient still remains statistically significantly far from 1, and the log-likelihood ratio test still favors the GCPC regression over the CIPC regression model.
Table 5: Estimated regression parameters (their standard errors appear within parentheses) for the GCPC regression model fitted to the wind data. On the right is the corrected version of the GCPC regression model.
Model GCPC (Tsagris and Alzeley,, 2025) GCPC
$\cos(y)$ $\sin(y)$ $\cos(y)$ $\sin(y)$
$\hat{\bm{\alpha}} -$0.042 (0.0013) -0.045 (0.0023) 0.164 (0.0012) 0.195 (0.0012)
$\hat{\bm{\beta}}$ 0.009 (0.0005) 0.010 (0.0009) -0.010 (0.0022) -0.012 (0.0021)
Loglik -331.846 -335.219
$\hat{\rho}$ 0.203 (0.039) 0.226 (0.055)
5 Conclusions
We derived the relationship between the GCPC and the CIPC (reparameterized wrapped Cauchy) distribution, provided a not so convenient formula for the mean resultant length, an alternative analytical formula to compute probabilities, and an analytical formula for the entropy of the GCPC. We further proposed two log-likelihood ratio tests for hypothesis testing with one or two location parameters. For the two location parameters testing case, the test is size correct when the data are wrongfully assumed to follow the CIPC distribution. Finally we corrected a mistake in the regression setting and proposed a faster altenative than the one proposed by Tsagris and Alzeley, (2025).

### 25 Mar 2026 v3

@@
 Abstract
-Tsagris and Alzeley, (2025) proposed the generalized circular projected Cauchy distribution, whose special case is the wrapped Cauchy distribution.
-In this paper we first derive the relationship with the wrapped Cauchy distribution, and then we attempt to characerize the distribution.
-We provide a non-analytical formula for the mean resultant length, and analytical form for the cumulative probability function and the the entropy.
-We propose a log-likelihood ratio test for the equality of two angular means, without assuming equality of thew concentration parameters, and briefly mention the addition of circular and simplicial predictors.
-Simulation studies illustrate a) the performance of the log-likelihood ratio test when one falsely assumes that the true distribution is the wrapped Cauchy distribution, and b) the empirical rate of covergence of the regression coefficients.
-Using a real data analysis example we correct a mistake in the regression setting.
-Keywords: Circular data; mean resultant length, hypothesis test
+Tsagris and Alzeley (2025) proposed the generalized circular projected Cauchy (GCPC) distribution, whose special case is the wrapped Cauchy distribution.
+In this paper we first derive the relationship with the wrapped Cauchy distribution, and then we attempt to characterize the distribution.
+We establish the conditions under which the distribution exhibits unimodality.
+We provide non-analytical formulas for the mean resultant length and the Kullback-Leibler divergence, and analytical form for the cumulative probability function and the entropy of the GCPC distribution.
+We propose log-likelihood ratio tests for one, or two location parameters without assuming equality of the concentration parameters.
+We revisit maximum likelihood estimation with and without predictors.
+In the regression setting we briefly mention the addition of circular and simplicial predictors.
+Simulation studies illustrate a) the performance of the log-likelihood ratio test when one falsely assumes that the true distribution is the wrapped Cauchy distribution, and b) the empirical rate of convergence of the regression coefficients.
+Using a real data analysis example we show how to avoid the log-likelihood being trapped in a local maximum and we correct a mistake in the regression setting.
+Keywords: Circular data; distribution, hypothesis test, maximum likelihood estimation
 MSC: 62H11, 62H10
@@
 Circular data, when $d=2$, lie on a circle.
-Circular data are encountered in various disciplines, such as political sciences (Gill and Hangartner,, 2010), criminology (Shirota and Gelfand,, 2017), biology (Landler et al.,, 2018), ecology (Horne et al.,, 2007), and astronomy (Soler et al.,, 2019), to name but a few.
-A large class of distributions has been proposed, see Tsagris and Alzeley, (2025) for a short list.
-In this paper we focus on the generalized circular projected Cauchy distribution (GCPC) proposed by Tsagris and Alzeley, (2025).
-Specifically we show the relationship with the wrapped Cauchy (WC) distribution and further propose a log-likelihood ratio test for the equality of two independent angular means.
-The next section briefly presents the GCPC distribution, derives the relationship with the WC distribution and proposes the log-likelihood ratio test.
-Then, simulation studies follow and the conclusions that close the paper.
+Circular data are encountered in various disciplines, such as political sciences (Gill and Hangartner, 2010), criminology (Shirota and Gelfand, 2017), biology (Landler et al., 2018), ecology (Horne et al., 2007), and astronomy (Soler et al., 2019), to name but a few.
+A large class of distributions has been proposed, see Tsagris and Alzeley (2025) for a short list.
+A classic distribution is the wrapped Cauchy distribution (Kent and Tyler, 1988), for which Tsagris and Alzeley (2025) showed to be a special case of the generalized projected Cauchy distribution (GCPC).
+The GCPC generalizes the WC by the introduction of an extra parameter that allows for anisotropy.
+The benefit of the GCPC distribution is that it provides a better fit to asymmetric and bimodal data.
+In this paper we focus on the GCPC distribution.
+Specifically we derive the relationship with the wrapped Cauchy (WC) distribution.
+We examine the conditions for unimodality and then provide an alternative formula for the cumulative probability function.
+We derive non-analytical formulas for its mean resultant length and the Kullback-Leibler divergence from the WC distribution, but derive analytical formula for its entropy.
+We further propose two log-likelihood ratio tests for equality of one and two location parameters, without assuming equal concentration parameters.
+We revisit the maximum likelihood estimation (MLE) and the regression modelling.
+For the MLE we show a problem that can trap the log-likelihood in a local maximum and show how to easily escape and reach the global maximum.
+We further correct a mistake of Tsagris and Alzeley (2025) and examine empirically the convergence rate of the regression coefficients.
+A real data analysis, with and without predictor variables illustrates the superior performance of the GCPC compared to the WC distribution.
+The next section briefly presents the GCPC distribution, and studies the distribution. Next, simulation studies and the real data example follow, with conclusions closing the paper.
 2 The GCPC distribution
@@
 $f({\bf x})=\frac{1}{2\pi|\boldsymbol{\Sigma}|^{1/2}}\left[1+\left({\bf x}-\boldsymbol{\mu}\right)^{\top}\boldsymbol{\Sigma}^{-1}\left({\bf x}-\boldsymbol{\mu}\right)\right]^{-3/2}.$ (2)
-By substituting (2) into (1) and evaluating the integral, Tsagris and Alzeley, (2025) derived the termed the Circular Projected Cauchy (CPC) distribution.
+By substituting (2) into (1) and evaluating the integral, Tsagris and Alzeley (2025) derived the circular projected Cauchy (CPC) distribution.
 $f({\bf y})$ $=$ $\int_{0}^{\infty}\frac{r}{2\pi|\boldsymbol{\Sigma}|^{1/2}}\left(1+r^{2}{\bf y}^{\top}\boldsymbol{\Sigma}^{-1}{\bf y}-2r{\bf y}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\mu}+\boldsymbol{\mu}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\mu}\right)^{-3/2}dr$ (3)
@@
 It is important to note that ${\bf y}\in\mathbb{S}^{1}$, while $\bm{\mu}\in\mathbb{R}^{2}$.
-(Tsagris and Alzeley,, 2025) relaxed this strict assumption by employing one of the conditions imposed in Paine et al., (2018), that is, $\boldsymbol{\Sigma\mu}=\boldsymbol{\mu}$, but not $|\boldsymbol{\Sigma}|=1$.
+(Tsagris and Alzeley, 2025) relaxed this strict assumption by employing one of the conditions imposed in Paine et al.
+(2018), that is, $\boldsymbol{\Sigma\mu}=\boldsymbol{\mu}$, but not $|\boldsymbol{\Sigma}|=1$.
 This condition implies that the one eigenvector $\boldsymbol{\xi}_{2}$ of $\boldsymbol{\Sigma}$ is the normalised location vector $\boldsymbol{\xi}_{2}=\left(\mu_{1},\mu_{2}\right)^{\top}/\gamma$, while the other eigenvector can be defined up to the sign as $\boldsymbol{\xi}_{1}=\left(-\mu_{2},\mu_{1}\right)^{\top}/\gamma$ or $\boldsymbol{\xi}_{1}=\left(\mu_{2},-\mu_{1}\right)^{\top}/\gamma$.
@@
 where $a=\gamma\cos\phi,\ b=\cos^{2}\phi+\frac{\sin^{2}\phi}{\lambda},\ \phi=\theta-\omega$, with $\lambda>0$, $\gamma\geq 0$ and $\omega\in[-\pi,\pi]$.
-The GCPC distribution exhibits symmetry with respect to $\omega$ such that $f(\theta-\omega)=f(\omega-\theta)$, and is thus symmetrical about $\omega+\pi$. The density of the GCPC may be re-written as
+We shall denote the eigenvalue, $\lambda$, of the covariance matrix $\bm{\Sigma}$ by anisotropy parameter, and the reason is explained in the next subsection.
+The GCPC distribution exhibits reflective symmetry with respect to $\omega$ only if $\lambda=1$, but its density function is even since $f(\theta-\omega)=f(\omega-\theta)$.
+The maximum value of the density occurs when $\theta=\omega$ and its value is $\left[2\pi\sqrt{\lambda}\left(\sqrt{\gamma^{2}+1}-\gamma\right)\right]^{-1}$.
+The density of the GCPC may be re-written as
 $f(\theta)=\frac{1}{2\pi\sqrt{\lambda}\sqrt{b}D}=\frac{1}{2\pi\sqrt{1+\left(\lambda-1\right)\cos^{2}{\phi}}\cdot D},$ (9)
@@
 $D=c\sqrt{b}-\gamma\cos\phi\ \text{and}\ c=\sqrt{\gamma^{2}+1}.$
-It is important to note that if $\lambda=1$, the GCPC distribution reduces to the circular independent projected Cauchy (CIPC) distribution (Tsagris et al.,, 2025)
+It is important to note that if the anisotropy parameter, $\lambda=1$, the GCPC distribution reduces to the circular independent projected Cauchy (CIPC) distribution (Tsagris et al., 2025)
 $f_{CIPC}({\bf\theta})=\dfrac{1}{2{\pi}\left(\sqrt{\gamma^{2}+1}-\gamma\cos{\left(\theta-\omega\right)}\right)},$ (10)
-which is the WC distribution (Mardia and Jupp,, 2000, pg. 51) with a different parameterization
+which is the WC distribution (Mardia and Jupp, 2000, pg. 51) with a different parameterization
 $f_{WC}(\theta)=\frac{1-\delta^{2}}{2\pi\left[1+\delta^{2}-2\delta\cos\left(\theta-\omega\right)\right]},$ (11)
-where $\gamma=\frac{2\delta}{1-\delta^{2}}$ or, conversely, $\delta=(\sqrt{\gamma^{2}+1}-1)/\gamma$ (Tsagris and Alzeley,, 2025).
+where $\gamma=\frac{2\delta}{1-\delta^{2}}$ or, conversely, $\delta=(\sqrt{\gamma^{2}+1}-1)/\gamma$ (Tsagris and Alzeley, 2025).
+As the name suggests, $\lambda$ controls the anisotropy of the GCPC, and if $\lambda=1$ we end up with an isotropic covariance matrix $\bm{\Sigma}$.
 2.1 Relationship with the CIPC distribution
@@
 Proof.
-The proof is straightfoward by application of the change-of-variables formula to obtain
-$f_{\Psi}(\psi)=f_{\Theta}(\theta)\cdot\left|\frac{d\theta}{d\psi}\right|=\frac{1}{2\pi\sqrt{\lambda}\,\sqrt{b}\,D}\cdot\sqrt{\lambda}\cdot b=\frac{1}{2\pi\!\left(\sqrt{\gamma^{2}+1}-\gamma\cos\psi\right)}.$
+The proof is straightforward by application of the change-of-variables formula.
+Without loss of generality, set $\omega=0$.
+Under the transformation $\psi=\arctan 2(\sin\varphi,\sqrt{\lambda}\cos\varphi)$ we can see that
+$\cos\psi=\frac{\cos\varphi}{\sqrt{b}},\qquad\sin\psi=\frac{\sin\varphi}{\sqrt{\lambda b}},$
+where $b=\cos^{2}\varphi+\sin^{2}\varphi/\lambda$.
+Differentiating $\tan\psi=\tan\varphi/\sqrt{\lambda}$ gives $|d\varphi/d\psi|=\sqrt{\lambda}\,b$.
+Applying the change-of-variables formula to $f_{\Phi}(\varphi)=(2\pi\sqrt{\lambda\,b}\,D)^{-1}$, with $D=c\sqrt{b}-\gamma\cos\varphi$ and $c=\sqrt{\gamma^{2}+1}$, we obtain:
+$f_{\Psi}(\psi)\;=\;f_{\Phi}(\varphi)\cdot\sqrt{\lambda}b\;=\;\frac{\sqrt{b}}{2\pi D}\;=\;\frac{\sqrt{b}}{2\pi(c\sqrt{b}-\gamma\sqrt{b}\cos\psi)}\;=\;\frac{1}{2\pi(\sqrt{\gamma^{2}+1}-\gamma\cos\psi)},$
+where in the last step we used $\cos\varphi=\sqrt{b}\cos\psi$. This is the $\mathrm{CIPC}(0,\gamma)\equiv\mathrm{WC}(0,\delta)$ density. ∎
+Based on this we can define the opposite transformation as follows:
+Lemma 2.1.
+[…]
+Proof.
+The proof is again straightforward and hence omitted. ∎
+2.2 Unimodality conditions for the GCPC distribution
+Tsagris and Alzeley (2025) observed that the density function of the GCPC (8) may be bimodal. They equated the derivative of the log-density to zero and obtained:
+$\sin(\theta-\omega)\left[2(\lambda-1)\sqrt{\gamma^{2}+1}\cos(\theta-\omega)\sqrt{\frac{(\lambda-1)\cos^{2}(\theta-\omega)+1}{\lambda}}-2(\lambda-1)\gamma\cos^{2}(\theta-\omega)-\gamma\right]$ $=$ $0.$ (19)
+This yields two cases: either $\sin(\theta-\omega)=0$ or the expression in brackets equals zero. Case 1: $\sin(\theta-\omega)=0$.
+$\theta-\omega=k\pi\implies\theta=\omega+k\pi,\ \ k\in\mathbb{Z}.$
+Case 2: The term inside the brackets equals zero. Let $u=\cos(\theta-\omega)$:
+$2(\lambda-1)\sqrt{\gamma^{2}+1}u\sqrt{\frac{(\lambda-1)u^{2}+1}{\lambda}}$ $=$ $\gamma\left(2(\lambda-1)u^{2}+1\right)$
+$4(\lambda-1)^{2}(\gamma^{2}+1)\,u^{2}\frac{(\lambda-1)u^{2}+1}{\lambda}$ $=$ $\gamma^{2}\!\left(2(\lambda-1)u^{2}+1\right)^{2}\ \ \text{(By squaring both sides)}.$
+Letting $t=u^{2}=\cos^{2}(\theta-\omega)$ and multiplying both sides with $\lambda$ yields:
+$4(\lambda-1)^{3}(\gamma^{2}+1)\,t^{2}+4(\lambda-1)^{2}(\gamma^{2}+1)t$ $=$ $4\lambda\gamma^{2}(\lambda-1)^{2}t^{2}+4\lambda\gamma^{2}(\lambda-1)+\lambda\gamma^{2}.$
+$4(\lambda-1)^{2}Mt^{2}+4(\lambda-1)Mt-\lambda\gamma^{2}$ $=$ $0,$
+where $M=\lambda-\gamma^{2}-1$. Hence,
+$t^{*}=\cos^{2}(\theta-\omega)=\frac{-M\pm\sqrt{M\left(M+\lambda\gamma^{2}\right)}}{2(\lambda-1)M}.$
+Finally,
+$\theta=\omega\pm\arccos\!\sqrt{t^{*}}.$
+Notes.
+  • If $M=0\implies\lambda=\gamma^{2}+1$, the quadratic degenerates and requires $\gamma=0$.
+  • The term inside the square term of $t^{*}$ expands as $M\left(M+\lambda\gamma^{2}\right)=\left(\lambda-\gamma^{2}-1\right)\left[\left(\lambda-1\right)\left(\gamma^{2}+1\right)\right]$.
+To ensure that this is non-negative two conditions apply:
+    • If $\lambda<1$ then $\lambda<\gamma^{2}+1$.
+    • If $\lambda>1$ then $\lambda>\gamma^{2}+1$.
+  • If $\lambda=1$, the term inside the brackets in (19) vanishes and we end up with Case 1 (unimodality of the distribution).
+  • If $\lambda\neq 1$, then the following four cases apply:
+    • Case A: $1<\lambda\leq\gamma^{2}+1$. The distribution is always unimodal since $M(\lambda-1)\leq 0$ gives complex roots, with no further conditions needed.
+    • Case B: $\lambda>\gamma^{2}+1$. The distribution is unimodal when $t^{*}_{+}>1$, i.e. when:
+      $(\gamma^{2}+1)(\lambda-1)>(\lambda-\gamma^{2}-1)(2\lambda-1)^{2}.$
+    • Case C: $\frac{1}{2}<\lambda<1$. The distribution is unimodal when $t^{*}_{-}>1$, i.e. when:
+      $(\gamma^{2}+1-\lambda)(1-2\lambda)^{2}>(\gamma^{2}+1)(1-\lambda).$
+    • Case D: $\lambda\leq\frac{1}{2}$. Here $t^{*}_{-}\in[0,1]$ always, so the distribution is never unimodal for $\lambda\leq 1/2$ regardless of $\gamma$; bimodal critical points always exist.
+The conditions stated in Cases A and D are straightforward for the bimodality.
+Cases B and C state that if $t^{*}$ falls outside the admissible region, the distribution is unimodal, where $t^{*}_{+}=\frac{M+\sqrt{M\left(M+\lambda\gamma^{2}\right)}}{2(\lambda-1)M}$ and $t^{*}_{-}=\frac{-M-\sqrt{M\left(M+\lambda\gamma^{2}\right)}}{2(\lambda-1)M}$.
+In those two cases, examination of bimodality requires some extra computations.
 2.3 Probabilities
-Following Tsagris and Alzeley, (2025), instead of the cumulative probability function we derive the probability included within a given interval $(a,b)$, and provide an alternative formula.
+Following Tsagris and Alzeley (2025), instead of the cumulative probability function we derive the probability included within a given interval $(a,b)$, and provide an alternative formula.
 $P(a\leq\theta\leq b)=\int_{a}^{b}f_{GCPC}(t)dt=\int_{-\pi}^{b}f_{GCPC}(t)dt-\int_{-\pi}^{a}f_{GCPC}(t)dt=F_{GCPC}(b)-F_{GCPC}(a).$
@@
 Based on this, one may compute the circular variance as $1-\rho$, and the circular standard deviation as $\left(-2\log{\rho}\right)^{1/2}$.
-Using Theorem 2.1 we may express this expectation as
-$\rho=\frac{1-\delta^{2}}{\pi}\int_{0}^{\pi}\frac{\cos\psi}{\sqrt{\cos^{2}\psi+\lambda\sin^{2}\psi}}\cdot\frac{1}{1+\delta^{2}-2\delta\cos\psi}\,d\psi.$
-After some algebra we end up with the following representation
-$\rho=\frac{2(1+\delta)}{\pi(1-\delta)\sqrt{\lambda}}\;\Pi\!\left(\left(\frac{1+\delta}{1-\delta}\right)^{2}\frac{1}{\lambda}-1\;\Bigg|\;1-\frac{1}{\lambda^{2}}\right),$
-where $\Pi(a|b)$ is the complete elliptic integral of the third kind (Ward,, 1960).
-We still see that there is no closed form solution for $\rho$, unless $\lambda=1$, and this applies to higher trigonometric moments.
+Theorem $2.$
+The mean resultant length of the GCPC distribution is given by
+$\rho=\frac{2(1+\delta)}{\pi(1-\delta)\sqrt{\lambda}}\;\Pi\left(\left(\frac{1+\delta}{1-\delta}\right)^{2}\frac{1}{\lambda}-1,1-\frac{1}{\lambda^{2}}\right),$
+where $\Pi\left(n,k\right)$ is the complete elliptic integral of the third kind (Ward, 1960).
+Proof.
+$E[\cos\theta]=\frac{1}{2\pi\sqrt{\lambda}}\int_{-\pi}^{\pi}\frac{\cos\theta}{\sqrt{b}D}\,d\theta,\qquad b=\cos^{2}\theta+\frac{\sin^{2}\theta}{\lambda},\quad D=\sqrt{\gamma^{2}+1}\,\sqrt{b}-\gamma\cos\theta.$
+Applying the change of variables $\psi=\arctan 2\!\left(\sin\theta,\,\sqrt{\lambda}\cos\theta\right)$ from Theorem 2.1, so that $f_{\mathrm{GCPC}}(\theta)\,d\theta=f_{WC}(\psi)\,d\psi$, we express $\cos\theta$ in terms of $\psi$:
+$\cos\psi=\frac{\cos\theta}{\sqrt{b}},\qquad\sin\psi=\frac{\sin\theta}{\sqrt{\lambda b}},\qquad b=\frac{1}{1+(\lambda-1)\sin^{2}\psi},\qquad\cos\theta=\frac{\cos\psi}{\sqrt{1+(\lambda-1)\sin^{2}\psi}}.$
+Substituting and using $f_{WC}(\psi)=\dfrac{1-\delta^{2}}{2\pi(1+\delta^{2}-2\delta\cos\psi)}$ we obtain:
+$E\left(\cos\theta\right)=\frac{1-\delta^{2}}{2\pi}\int_{-\pi}^{\pi}\frac{\cos\psi}{\left(1+\delta^{2}-2\delta\cos\psi\right)\sqrt{1+(\lambda-1)\sin^{2}\psi}}\,d\psi.$
+Since the integrand is even in $\psi$, doubling the integral over $0,\pi]$ and reducing via the substitution $u=\pi/2-\psi$ yields:
+$E\left(\cos\theta\right)=\frac{2(1+\delta)}{\pi(1-\delta)\sqrt{\lambda}}\;\Pi\left(\left(\frac{1+\delta}{1-\delta}\right)^{\!2}\left(\frac{1}{\lambda}-1\right),\sqrt{1-\frac{1}{\lambda^{2}}}\right),$
+confirming the non-analytical nature of $\rho=E[\cos\theta]$ for $\lambda\neq 1$. ∎
+It remains the case that there is no analytical formula for $\rho$, unless $\lambda=1$, and this applies to higher trigonometric moments.
 Figure 1(a) visualizes the values of $\rho$ for a grid of values of $\gamma$ and $\lambda$.
@@
 The entropy of the GCPC distribution is given by
-$H_{GCPC}(\omega,\gamma,\lambda)=\log{\left\{\frac{8\pi\sqrt{\lambda}\,(1-\delta^{2})}{\bigl[\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}\bigr]^{2}}\right\}},$
-where $\delta$ was defined earlier.
+$H_{GCPC}(\omega,\gamma,\lambda)=\log{\left\{\frac{8\pi\sqrt{\lambda}\,(1-\delta^{2})}{\bigl[\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}\bigr]^{2}}\right\}}.$
 Proof.
@@
 Figure 1(b) visualizes the values of $H$ for a grid of values of $\gamma$ and $\lambda$.
-We observe that, in cotrast to $\rho$, $H$ increases with increasing $\lambda$ and decreases with increasing $\gamma$ values.
+We observe that, in contrast to $\rho$, $H$ increases with increasing $\lambda$ and decreases with increasing $\gamma$ values.
 Refer to caption Refer to caption
@@
 where the expectation is taken with respect to $\psi\sim WC(0,\delta)$.
+2.7 Maximum likelihood estimation
+Tsagris and Alzeley (2025) performed maximum likelihood estimation (MLE) using the Euclidean representation of the density of the GCPC (7).
+We observed that this approach is sub-optimal and can yield a local instead of the global maximum, when the distribution is bimodal.
+We emphasize though that if the distribution is unimodal their approach is valid.
+Despite their efforts to ensure the maximum, via a clever implementation, we will show in the real data analysis, that the use of the log-likelihood parametrized using 7 does not lead to the global maximum, since the log-likelihood can have four local maxima [^1] Tsagris and Alzeley (2025) mentioned that the solution to Eq.
+(19) has four roots., when the distribution is bimodal.
+Our solution is to plot the data first and observe whether the distribution is unimodal or bimodal and then decide on which density parametrization to employ.
 2.8 Hypothesis testing and confidence intervals
@@
 In order to test whether the location parameter equals some pre-specified value we will employ the log-likelihood ratio test.
-Under the $H_{0}$, $\omega=\omega_{0}$, and so we need to maximize the restricted log-likelihood, $\ell_{0}$, with respect to the other two parameters, $\gamma$ and $\lambda$, and estimate their values, $\tilde{gamma}$ and $\tilde{\lambda}$.
+Under the $H_{0}$, $\omega=\omega_{0}$, and so we need to maximize the restricted log-likelihood, $\ell_{0}$, with respect to the other two parameters, $\gamma$ and $\lambda$, and estimate their values, $\tilde{\gamma}$ and $\tilde{\lambda}$.
 Under the $H_{1}$, $\omega\neq\omega_{0}$, and we maximize the unrestricted log-likelihood, $\ell_{1}$, to obtain $\hat{\omega}$, $\hat{\gamma}$ and $\hat{\lambda}$.
 Under regularity conditions, $\Lambda=2\left[\ell_{1}\left(\hat{\omega},\hat{\omega},\hat{\lambda}\right)-\ell_{0}\left(\omega_{0},\tilde{\gamma},\tilde{\lambda}\right)\right]\mathrel{\dot{\sim}}\chi^{2}_{1}$.
-2.6.2 Confidence interval for a location parameter
-Using the log-likelihood ratio test, we can construct asymptotic confidence intervals for the true location parameter $\omega$ of a sample by searching for the pair of values for which it holds that (Cox and Hinkley,, 1979): $\ell\left(\omega,\hat{\gamma},\hat{\lambda}\right)>\ell\left(\hat{\omega},\hat{\gamma},\hat{\lambda}\right)-\frac{\chi^{2}_{1}}{2}$.
+2.8.2 Confidence interval for the true location parameter
+Using the log-likelihood ratio test, we can construct asymptotic confidence intervals for the true location parameter $\omega$ of a sample by searching for the pair of values for which it holds that (Cox and Hinkley, 1979): $\ell\left(\omega,\hat{\gamma},\hat{\lambda}\right)>\ell\left(\hat{\omega},\hat{\gamma},\hat{\lambda}\right)-\frac{\chi^{2}_{1}}{2}$.
 2.8.3 Equality of two location parameters
-To test the equality of two location parameters, without assuming equality of the parameters $\gamma$ and $\lambda$, we will perform a log-likelihood ratio test [^1] The function to perform the log-likelihood ratio test is available in the R package Directional (Tsagris et al.,, 2026).
-following (Tsagris et al.,, 2025).
-Assume we have circular observations from two independent samples, $(\theta_{11},\ldots,\theta_{1{n_{1}}})$ and $(\theta_{11},\ldots,\theta_{1{n_{2}}})$, where $n_{1}$ and $n_{2}$ denote the sample sizes of the two samples.
+To test the equality of two location parameters, without assuming equality of the parameters $\gamma$ and $\lambda$, we will perform a log-likelihood ratio test [^2] All the relevant functions are available in the R package Directional (Tsagris et al., 2026).
+following (Tsagris et al., 2025).
+Assume we have circular observations from two independent samples, $(\theta_{11},\ldots,\theta_{1{n_{1}}})$ and $(\theta_{21},\ldots,\theta_{2{n_{2}}})$, where $n_{1}$ and $n_{2}$ denote the sample sizes of the two samples.
 Under $H_{0}$, $\omega_{1}=\omega_{2}=\tilde{\omega}$, and by using Eq. (8), the log-likelihood is written as
@@
 $-\sum_{i=1}^{n_{1}}\log{\left(\tilde{b}_{1i}\sqrt{\tilde{\gamma}_{1}^{2}+1}-\tilde{a}_{1i}\sqrt{\tilde{b}_{1i}}\right)}-\sum_{i=1}^{n_{2}}\log{\left(\tilde{b}_{2i}\sqrt{\tilde{\gamma}_{2}^{2}+1}-\tilde{a}_{2i}\sqrt{\tilde{b}_{2i}}\right)},$
-where $\tilde{a}_{ji}=e^{\tilde{\gamma}_{j}}\cos{\left(\theta_{ji}-\tilde{\omega}\right)}$ and $\tilde{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\tilde{\omega}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\tilde{\omega}\right)}{\tilde{\lambda}_{j}}$, for $j=1,2$.
+where $\tilde{a}_{ji}=\tilde{\gamma}_{j}\cos{\left(\theta_{ji}-\tilde{\omega}\right)}$ and $\tilde{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\tilde{\omega}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\tilde{\omega}\right)}{\tilde{\lambda}_{j}}$, for $j=1,2$.
 Under $H_{1}$, $\omega_{1}\neq\omega_{2}$, and the log-likelihood is written as
@@
 $-\sum_{i=1}^{n_{1}}\log{\left(\hat{b}_{1i}\sqrt{\gamma_{1}^{2}+1}-\hat{a}_{1i}\sqrt{\hat{b}_{1i}}\right)}-\sum_{i=1}^{n_{2}}\log{\left(\hat{b}_{2i}\sqrt{\hat{\gamma}_{2}^{2}+1}-\hat{a}_{2i}\sqrt{\hat{b}_{2i}}\right)},$
-where $\hat{a}_{ji}=e^{\hat{\gamma}_{j}}\cos{\left(\theta_{ji}-\hat{\omega}_{j}\right)}$ and $\hat{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\hat{\omega}_{j}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\hat{\omega}_{j}\right)}{\hat{\lambda}_{j}}$, for $j=1,2$.
+where $\hat{a}_{ji}=\hat{\gamma}_{j}\cos{\left(\theta_{ji}-\hat{\omega}_{j}\right)}$ and $\hat{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\hat{\omega}_{j}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\hat{\omega}_{j}\right)}{\hat{\lambda}_{j}}$, for $j=1,2$.
 Under regularity conditions, $\Lambda=2\left[\ell_{1}\left(\hat{\omega},\hat{\lambda}_{1},\hat{\lambda}_{2},\hat{\gamma}_{1},\hat{\gamma}_{2}\right)-\ell_{0}\left(\tilde{\omega},\tilde{\lambda}_{1},\tilde{\lambda}_{2},\tilde{\gamma}_{1},\tilde{\gamma}_{2}\right)\right]\mathrel{\dot{\sim}}\chi^{2}_{1}$.
 2.9 Regression modelling revisited
-Tsagris and Alzeley, (2025) defined the GCPC regression by considering the bivariate representation (Euclidean coordinates) of the circular data as opposed to their univariate nature (Kato et al.,, 2008).
-This is akin to the approach employed in the SPML model, as detailed by Presnell et al., (1998) and allows for varying concentration parameter ($\lambda$).
+Tsagris and Alzeley (2025) defined the GCPC regression by considering the bivariate representation (Euclidean coordinates) of the circular data as opposed to their univariate nature (Kato et al., 2008).
+This is akin to the approach employed in the SPML model, as detailed by Presnell et al.
+(1998) and allows for varying concentration parameter ($\lambda$).
 The log-likelihood of the GCPC regression model is written as
@@
 where
-$\bm{y}_{i}^{\top}\bm{\Sigma}^{-1}_{i}\left(\bm{B}\right)\bm{y}_{i}$ $=$ $\xi_{1i}\left(\bm{B}\right)\xi_{1i}\left(\bm{B}\right)^{\top}/\lambda+\xi_{2i}\left(\bm{B}\right)\xi_{2i}\left(\bm{B}\right)^{\top}$
-$=$ $y_{1i}^{2}\left(\xi_{1i}^{2}/\lambda+\xi_{2i}^{2}\right)+y_{2i}^{2}\left(\xi_{2i}^{2}/\lambda+\xi_{1i}^{2}\right)+2y_{1i}y_{2i}\xi_{1i}\xi_{2i}\left(1-1/\lambda\right)$
+$\bm{y}_{i}^{\top}\bm{\Sigma}^{-1}_{i}\left(\bm{B}\right)\bm{y}_{i}$ $=$ $\bm{y}_{i}^{\top}\left[\xi_{1i}\left(\bm{B}\right)\xi_{1i}\left(\bm{B}\right)^{\top}/\lambda+\xi_{2i}\left(\bm{B}\right)\xi_{2i}\left(\bm{B}\right)^{\top}\right]\bm{y}_{i}$
+$=$ $y_{1i}^{2}\left(\xi_{2i}^{2}/\lambda+\xi_{1i}^{2}\right)+y_{2i}^{2}\left(\xi_{1i}^{2}/\lambda+\xi_{2i}^{2}\right)+2y_{1i}y_{2i}\xi_{1i}\xi_{2i}\left(1-1/\lambda\right)$
+$=$ $\left(y_{1i}\xi_{2i}-y_{2i}\xi_{1i}\right)^{2}/\lambda+\left(y_{1i}\xi_{1i}+y_{2i}\xi_{2i}\right)^{2}.$
 and
 $\xi_{ji}=\frac{\bm{\beta}_{j}^{\top}x_{i}}{\|\bm{\beta}_{j}^{\top}x_{i}\|}=\frac{\mu_{ji}}{\|\bm{\mu}_{i}\|}=\frac{\mu_{ji}}{\gamma_{i}}\ \ \text{for}\ \ j=1,2,$ (22)
-where $\bm{x}_{i}$ denoting the $i$-th observation of the covariate vector.
-To maximize the log-likelihood of the GCPC regression model (20) suggested to use multiple starting points, but we propose a faster option.
-We begin with initial regression coefficients produced by the SPML regression model of Presnell et al., (1998) and then estimate the $\lambda$ parameter using Brent’s algorithm [^2] This is available in R via the buit-in optimize() function.
-(Brent,, 2013).
-Then use these as starting points in R’s built-in optim() function.
-Tsagris and Alzeley, (2025) did not consider the case of circular-circular regreression, where a covariate $X$ is circular as well.
-But, this case is easily adopted in the above case scenario by transforming the circular covariate to Euclidean by using $\bm{z}_{i}=\left(\cos{x_{i}},\sin{x_{i}}\right)^{\top}$.
+with $\bm{x}_{i}$ denoting the $i$-th observation of the covariate vector.
+To maximize the log-likelihood of the GCPC regression model (21), Tsagris and Alzeley (2025) suggested the use of multiple starting values, however we propose a computationally more efficient approach.
+We begin with initial regression coefficients produced by the SPML regression model of Presnell et al.
+(1998) and then estimate the $\lambda$ parameter using Brent’s algorithm [^3] This is available in R via the built-in optimize() function.
+(Brent, 2013).
+These estimates are subsequently used as starting values in R’s built-in optim() function.
+Tsagris and Alzeley (2025) did not consider the case of circular-circular regression, where a covariate $X$ is circular.
+But, this case is easily accommodated in the above case scenario by transforming the circular covariate to Euclidean by using $\bm{z}_{i}=\left(\cos{x_{i}},\sin{x_{i}}\right)^{\top}$.
 In case of multiple circular covariates, all of them are transformed into their Euclidean coordinates and the same link function (22) is used.
-Simplicial predictors (compositional data) are straighforward to add by using two options.
-The first is to apply the additive log-ratio transformation (Aitchison,, 2003), so that the regression coefficients sum to zero and have a nice (derivative based) interpretation.
-In case of zero values present the logarithmic transformation breaks down, so the alternative approach is to transform them via the $\alpha$–transformation (Tsagris et al.,, 2011).
-The second approach bypasses the problem of zero values but lacks interpretation of the estimated regression coefficients and requires to choose the value of the $\alpha$ parameter.
-The choice of $\alpha$ may be overcome by choosing a small value of $\alpha$ (say 0.01) to resemble the isometric log-ratio transformation (Egozcue et al.,, 2003).
-An extra extension would be to link the $\lambda$ parameter to the covariates as well, but that would increase the computational cost.
-Another evident approach to make the model non-linear would be to use splines in the covariates as well.
+Simplicial predictors (compositional data) are straightforward to add by using two options.
+The first is to apply the additive log-ratio transformation (Aitchison, 2003), so that the regression coefficients sum to zero and have a meaningful (derivative based) interpretation.
+In case of zero values present the logarithmic transformation breaks down, so the alternative approach is to transform them via the $\alpha$–transformation (Tsagris et al., 2011).
+The second approach bypasses the problem of zero values but lacks interpretation of the estimated regression coefficients and requires one to select the value of the $\alpha$ parameter.
+The choice of $\alpha$ may be overcome by choosing a small value of $\alpha$ (for instance, $\alpha=0.01$) to resemble the isometric log-ratio transformation (Egozcue et al., 2003).
+A further extension consists of linking the $\lambda$ parameter to the covariates, but that would increase the computational cost.
+A complementary approach to introduce non-linearity consists of incorporating splines for the covariate effects.
 3 Simulation studies
-We performed simulation studies to examine the performance of the two sample means test and also the effect of the $\lambda$ parameter on the performance of the log-likelihood test.
+We performed simulation studies to examine the performance of the test for one and two location parameters and to assess the effect of misspecifying $\lambda$ on the size of the log-likelihood ratio test.
+We did not perform simulation studies for the asymptotic confidence interval for the location parameter since this is derived from the log-likelihood ratio test.
 3.1 One location parameter
@@
 For each sample size we computed the type I error (assuming $\alpha=0.05$) based on 1,000 replicates.
-Table 1 presents the results, where we can see that the test slightly overestimates the nominal type I error, but not significantly.
+Table 1 presents the results, where we can observe that the test slightly overestimates the nominal type I error, though the deviation is not substantial.
 Table 1: Estimated type I error of the log-likelihood ratio test for one location parameter assuming the GCPC distribution.
@@
 We compared the test when both samples come from the GCPC distribution and when assuming that they come from the CIPC distribution ($\lambda=1$).
-Six pairs of unequal sample sizes were used, where the true angular mean was $\omega=2$, but the $\gamma$ and $\lambda$ parameters differed.
+Six pairs of unequal sample sizes were used, where the true location parameters were equal to $\omega=2$ for both populations, but the $\gamma$ and $\lambda$ parameters differed between the two populations.
 The parameters for the smaller sample were equal to $\gamma=4$ and $\lambda=1$, while for the larger sample they were equal to $\gamma=2$ and $\lambda=3$.
@@
 $\left(y_{1},\ y_{2}\right)=\left(1,\ \cos{u},\ \sin{u}\right)\left(\begin{array}[]{cc}\beta_{01}&\beta_{02}\\ \beta_{11}&\beta_{12}\\ \beta_{21}&\beta_{22}\end{array}\right),$
-where $\bm{u}$ was generated from a von Mises distribution with mean equal to 2 and concentration parameter equal to 5. For the continuous predictor, the model was
+where $\bm{u}$ was generated from a von Mises distribution with location parameter $\omega=2$ and concentration parameter $\gamma=5$. For the continuous predictor, the model was
 $\left(y_{1},\ y_{2}\right)=\left(1,\ x\right)\left(\begin{array}[]{cc}\beta_{01}&\beta_{02}\\ \beta_{11}&\beta_{12}\\ \end{array}\right),$
@@
 The first two columns correspond to the coefficients of the model with the circular predictor, the second two to the model with the continuous predictor and the last two to the model with both of them.
-In the last case $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$ refer to the circular predictor and $\bm{\beta}_{3}$ to the continous predictor.
+In the last case $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$ refer to the circular predictor and $\bm{\beta}_{3}$ to the continuous predictor.
 Predictor Circular Continuous Circular and continuous
@@
 (e) $\lambda=0.5$ (f) $\lambda=5$
-Figure 2: Average squared Frobenius norm $\bar{F}^{2}$ versus sample size, both in log-scale for the circular, the continuous and the combination of both.
+Figure 2: Average squared Frobenius norm $\bar{F}^{2}$ versus sample size, both in log-scale for the circular predictor, continuous predictor and the combination of both.
 4 Real data analysis
-We used the same dataset used in Tsagris and Alzeley, (2025), the speed.wind2 data that is available in the R package NPCirc (Oliveira et al.,, 2014).
+We used the same dataset used in Tsagris and Alzeley (2025), the speed.wind2 data that is available in the R package NPCirc (Oliveira et al., 2014).
 This dataset consists of 199 hourly observations of wind direction and wind speed in winter season (from November to February) from 2003 until 2012 in the Atlantic coast of Galicia (Spain).
-4.1 Maximum likelihood estimation
+4.1 MLE
+Table 4 presents the MLE of the CIPC and the GCPC distributions fitted to the speed direction data.
+We highlight that the MLE using Eq.
+(7), Tsagris and Alzeley (2025) was trapped in one of the three local maxima, while the MLE using (8) obtained the global maximum.
+Notice that the estimated $\lambda$ falls within Case D, $\lambda\leq\frac{1}{2}$, and hence the distribution is bimodal, which is also evident from Figures 3(b) and 3(c).
 Table 4 presents the estimated parameters of the CIPC and GCPC models.
-Application of the log-likelihood ratio test to discriminate between the two models (Tsagris and Alzeley,, 2025) clearly favors the GCPC model.
-Figure 3(a) visualizes the data, it is the circular plot with the mean directions showed with arrows.
-Evidently the data look bimodal and the CIPC has chosen one of the two modes, whereas the GCPC has chosen a direction that ”cuts” the data into two.
-The density plot on Figure 3(b) shows that the GCPC density has a similar pattern to the one produced by the kernel density estimate, in contrast to the CIPC distribution that has failed to fit the data.
+Application of the log-likelihood ratio test to discriminate between the two models (Tsagris and Alzeley, 2025) clearly favors the GCPC model.
+Figure 3(a) visualizes the data, it is the circular plot with the location parameter shown with arrows.
+The density plot on Figure 3(b) shows that the GCPC density has a similar pattern to the one produced by the kernel density estimate, in contrast to the CIPC distribution that does not adequately capture the distributional features of the data.
 This is more evident in the circular density plot (Figure 3(c)).
 The CIPC distribution has a circular shape, whereas the GCPC distribution has an elliptical shape, capturing the shape of the data more accurately.
-Finally, Figure 3(d) visualizes the 95% confidence interval for the mean direction.
-Table 4: Estimated parameters of the CIPC and GCPC models applied to the wind direction data.
+Figure 3(d) visualizes the 95% confidence interval for the location parameter.
+Figure 4 visualizes the profile log-likelihood of the location parameter $\omega$.
+The log-likelihood contains four modes, three of them correspond to local maxima and one is the global maximum.
+Table 4: Estimated parameters of the CIPC and GCPC models applied to the wind direction data. The GCPC* indicates the MLE using the polar coordinates parametrized GCPC density (8).
 Model $\hat{\omega}$ $\hat{\gamma}$ $\hat{\lambda}$ Log-likelihood
@@
 GCPC 5.587 0.050 4.21 -337.739
+GCPC* 0.873 0.238 0.155 -336.682
 Refer to caption Refer to caption
@@
 Refer to caption Refer to caption
-(c) Circular density plot (d) 95% confidence interval for the mean direction
+(c) Circular density plot (d) 95% confidence interval for the true location parameter
 Figure 3: (a) Circular plot of the speed direction data.
@@
 (c) Circular density plot of the two fitted distributions.
-The green line is the estimated mean direction based on the CIPC, whereas the blue lines is the estimated mean direction based on the GCPC model.
+The green line is the estimated location parameter based on the CIPC, whereas the blue line is the estimated location parameter based on the GCPC model.
+Figure 4: Profile log-likelihood of the location parameter $\omega$ for the GCPC distribution using the speed direction data.
+The white vertical line corresponds to the estimated location parameter that maximizes the log-likelihood.The periodicity is equal to $\pi/2$.
 4.2 Regression
-We spotted a mistake in the regression modelling of Tsagris and Alzeley, (2025), where the continuous predictor is the speed of the wind.
-Table 5 shows the regression coefficients computed in Tsagris and Alzeley, (2025) and the correct estimates.
-Note that this time the sign of the slope coefficients matches that of the CIPC and of the spherically projected multivariate linear regression models models.
+We identified a mistake in the regression modelling of Tsagris and Alzeley (2025), where the continuous predictor is the speed of the wind.
+Table 5 shows the regression coefficients computed in Tsagris and Alzeley (2025) and the correct estimates.
+Note that this time the sign of the slope coefficients matches that of the CIPC and of the spherically projected multivariate linear regression models.
 The $\lambda$ coefficient still remains statistically significantly far from 1, and the log-likelihood ratio test still favors the GCPC regression over the CIPC regression model.
@@
 On the right is the corrected version of the GCPC regression model.
-Model GCPC (Tsagris and Alzeley,, 2025) GCPC
+Model GCPC (Tsagris and Alzeley, 2025) GCPC
 $\cos(y)$ $\sin(y)$ $\cos(y)$ $\sin(y)$
@@
 5 Conclusions
-We derived the relationship between the GCPC and the CIPC (reparameterized wrapped Cauchy) distribution, provided a not so convenient formula for the mean resultant length, an alternative analytical formula to compute probabilities, and an analytical formula for the entropy of the GCPC.
+We derived the relationship between the GCPC and the CIPC (reparameterized WC) distribution, provided non-analytical formulas for the mean resultant length and the Kullback-Leibler divergence, an alternative analytical formula to compute probabilities, and an analytical formula for the entropy of the GCPC.
+We also showed the conditions under which the GCPC distribution is unimodal.
 We further proposed two log-likelihood ratio tests for hypothesis testing with one or two location parameters.
 For the two location parameters testing case, the test is size correct when the data are wrongfully assumed to follow the CIPC distribution.
-Finally we corrected a mistake in the regression setting and proposed a faster altenative than the one proposed by Tsagris and Alzeley, (2025).
+We observed that in the case of a bimodal distribution there is danger the MLE to be trapped in a local maximum and showed how to escape this trap.
+Finally, we corrected a mistake in the regression setting and proposed a computationally more efficient alternative than the one proposed by Tsagris and Alzeley (2025).